OpenStudy (anonymous):

A colony of bacteria is growing exponentially according to the function below, where t is in hours. How many hours will it take for the population to grow to 200? B(t) = 4 • e^0.8(t)

OpenStudy (nikvist):

\[B(t)=4\cdot\exp(0.8\cdot t)=200\]\[\exp(0.8\cdot t)=50\]\[t=1.25\cdot\ln{50}\approx4.89h\approx4h\,\,\,\,53min\]